# biblio.bib Resolution Limiting Factors in Low-Energy Cascade Zenith Reconstruction with the IceCube Upgrade

### v1

@@ 5 Results
The resolutions are shown in Fig. 9. The impact of module noise (solid red) on the overall resolution is minimal. This result is expected, given the low dark noise rate in mDOM PMTs resulting in a high signal purity, as discussed in Section 3. In the absence of scattering effects, the degradation of event information by photon direction smearing by PMTs (dashed green) is also negligible and comparable to module noise. This is because, despite the smearing of the angular distribution of photons caused by the PMTs, the charge distribution still enables an almost perfect reconstruction of the shower axis direction in the absence of a hadronic shower and photon scattering. The degradation in resolution due to the spread of shower particles (dashed-dotted orange) is significantly higher, making it a major factor in limiting event information. The resolution deterioration caused by in-ice photon scattering (dotted blue) is even greater than that from the shower spread highlighting that in-ice scattering is the primary factor restricting the physics information in events.
@@
The individual impacts of the information loss processes on reconstruction resolutions are analyzed in Fig. 9. Since these processes degrade reconstruction resolution, we evaluate the improvement in resolutions when any one of the resolution-limiting processes is disabled while keeping all others active in the simulation. This approach is the reverse of the earlier study, which analyzed the contribution of each limiting process to the overall resolution. The results in reconstruction resolutions are shown in Fig. 10. When the effect of PMT smearing of incident photon directions (dashed green) is removed, the degradation remains negligible indicating that module resolutions have a minimal impact on overall performance. An important implication of these results is that enhancing module resolutions by increasing the number of PMTs within a module would not significantly improve reconstruction performance. Disabling the effect of shower spread (dashed-dotted orange) yields an approximate 30% improvement in resolutions over the benchmark. The most pronounced improvement, around 50% over the benchmark, occurs when in-ice photon scattering (dotted blue) is excluded, indicating its dominant influence in limiting resolutions. While the former is inherent to shower formation, the latter could be reduced in a different detection medium with reduced scattering, such as water [P-ONE]. These results confirm the conclusions from earlier discussions regarding the contributions of resolution-limiting processes.
@@
The resolutions representing the combined degradation effect on reconstruction from all of the included information loss processes, were derived in Sec. 4.3 and have been shown as a reference line (black dotted) in Fig. 11. To determine the loss in directional resolutions due to PDF modeling errors in this case, a toy simulation is constructed. The opening angles and residual timings of hit photons from the PPC simulation are replaced with values sampled from the averaged integrated PDF. The observed values for each photon hit are replaced, ensuring that the integrated PDF and the per-module PDFs replicate the vertex-averaged photon PDFs, shown in the middle panel of Fig. 4. Since the observables, and , are randomly sampled from the averaged PDF, the simulation represents an unphysical scenario where the amount of scattering a photon goes through is distance independent. This means that a module close to an event might register larger timestamps and more scattered photons than a module much further away. However, for this study, this toy simulation simply demonstrates the reconstruction performance when the underlying angular and timing distributions are the same as the expected ones.
@@ 6 Summary and Conclusions
In ice-based neutrino detectors, inhomogeneities and anisotropies within the medium pose challenges (Section 3) for accurate shower reconstruction since they lead to increasing complexities in modeling the photon propagation in the ice. Moreover, new expected photon angular and timing distributions have to be generated for every new hypothesis which is challenging in likelihood reconstruction methods.
@@
These challenges have been circumvented by using homogeneous ice (Section 3) ensuring the obtained photon distributions are translationally and rotationally invariant. To bypass the issue of direction-dependent photon distributions in homogeneous ice due to the local detector geometry around the event, a randomized geometry is constructed from which averaged integrated PDFs are generated (Section 4). These PDFs are independent of incoming directions and are used for all simulated events. An extended unbinned likelihood reconstruction is employed based on the direction and timing information of individual photons along with the per-module charge information, simulated via the Photon Propagation Code (Section 4).
@@
We have derived the individual contributions from the processes limiting the physics information, which subsequently impose limitations on the achievable resolutions. We have demonstrated (Fig. 9) that in-ice photon scattering and transverse momentum spread of shower particles (Section 3) are the dominant factors in limiting physics information. These contributions dominate over limitations arising from the directional resolution of incoming photons from PMTs in modules and module noise, both exhibiting minimal impacts on the reconstruction degradation.
@@
The achievable reconstruction resolutions in ice are derived using integrated averaged PDFs and a toy simulation further indicates a 20% improvement in resolutions when angular and timing profiles of detected photons for every module are perfectly modeled instead of using an integrated distribution over all modules. These results highlight the scope for potential improvement in the performance of the state-of-the-art ML-based reconstruction algorithms on IceCube Upgrade simulations. Given the significant computational cost of training these algorithms, which increases with network complexity and benefits from larger training samples, these results provide key insights into optimizing resource allocation to enhance low-energy reconstruction methods.
@@
Although the simulations and subsequent reconstruction in this paper have been performed for ice, the analysis is applicable to water-based Cherenkov detectors like ORCA. This is because seawater and ice are predominantly H20 leading to similar neutrino-nucleon cross-sections and subsequent particle propagation and overall shower profiles [ORCA_shower]. Therefore, it is expected that the limits on directional resolution by the shower spread will apply equally well in seawater. The proposed detector layout in ORCA has multi-PMT optical modules similar to mDOMs which are deployed on vertical strings. This detector architecture is expected to record photon distributions comparable to those in this study, though they will be shifted due to the different refractive index of water. Due to the lower scattering coefficient of water, photon scattering would have a smaller effect on limiting resolutions in water than in ice.

### v2

@@ 5.1 Limitations in resolutions from individual information loss processes
The reconstruction resolutions on the zenith angle as a function of photon hit counts are shown in Fig. 9. The impact of module noise (solid red) on the overall resolution is minimal. This result is expected, given the low dark noise rate in mDOM PMTs resulting in a high signal purity, as discussed in Section 3. In the absence of scattering effects, the degradation of event information by photon direction smearing by PMTs (dashed green) is also negligible and comparable to module noise. This is because, despite the smearing of the angular distribution of photons caused by the PMTs, the charge distribution still enables an almost perfect reconstruction of the shower axis direction in the absence of a hadronic shower and photon scattering. The degradation in resolution due to the spread of shower particles (dashed-dotted burnt orange) is significantly higher, making it a major factor in limiting event information. The resolution deterioration caused by in-ice photon scattering (dotted blue) is even greater than that from the shower spread highlighting that in-ice scattering is the primary factor restricting the physics information in events.
@@
The individual impacts of the information loss processes on reconstruction resolutions are analyzed in Fig. 9. Since these processes degrade reconstruction resolution, we evaluate the improvement in resolutions when any one of the resolution-limiting processes is disabled while keeping all others active in the simulation. This approach is the reverse of the earlier study, which analyzed the contribution of each limiting process to the overall resolution. The results in reconstruction resolutions are shown in Fig. 10. When the effect of PMT smearing of incident photon directions (dashed green) is removed, the degradation remains negligible indicating that module resolutions have a minimal impact on overall performance. An important implication of these results is that improving on the directional accuracy of the hit photons does not contribute towards further improving our reconstruction. While we benefit from an increase in hit counts due to the larger effective area of the module with more PMTs, this is not relevant to the current study, as resolutions are shown as a function of photon hits rather than energies.
Disabling the effect of shower spread (dashed-dotted burnt orange) yields an approximate 30% improvement in resolutions over the benchmark. The most pronounced improvement, around 50% over the benchmark, occurs when in-ice photon scattering (dotted blue) is excluded, indicating its dominant influence in limiting resolutions. While the former is inherent to shower formation, the latter could be reduced in a different detection medium with reduced scattering, such as water [P-ONE]. The results confirm the conclusions from earlier discussions regarding the contributions of resolution-limiting processes. These systematics in IceCube dominate over the kinematic opening angle (Section 3; first paragraph) for NC and CC cascade events.
5.2 Achievable resolutions
@@
The resolutions representing the combined degradation effect on reconstruction from all of the included information loss processes, were derived in Sec. 4.3 and have been shown as a reference line (black dotted) in Fig. 11. To determine the loss in directional resolutions due to PDF modeling errors in this case, a toy simulation is constructed. The angle () and residual timings t of hit photons from the PPC simulation are replaced with values sampled from the averaged integrated PDF. The observed values for each photon hit are replaced, ensuring that the integrated PDF and the per-module PDFs replicate the vertex-averaged photon PDFs, shown in the middle panel of Fig. 4. Since the observables, and , are randomly sampled from the averaged PDF, the simulation represents an unphysical scenario where the amount of scattering a photon goes through is distance independent. This means that a module close to an event might register larger timestamps and more scattered photons than a module much further away. However, for this study, this toy simulation simply demonstrates the reconstruction performance when the underlying angular and timing distributions are the same as the expected ones.
@@ 6 Summary and Conclusions
In ice-based neutrino detectors, inhomogeneities and anisotropies within the medium pose challenges (Section 3) for accurate shower reconstruction since they complicate the modeling of photon propagation through ice. To obtain true likelihood profiles, it is necessary to generate expected angular and timing distributions of detected photons for each direction and vertex position hypothesis, which is computationally impractical for likelihood reconstruction methods.
@@
These challenges are circumvented by using homogeneous ice (Section 3) which ensures that the obtained photon distributions are translationally and rotationally invariant. To bypass the issue of direction-dependent photon distributions in homogeneous ice due to the local detector geometry around the event, a randomized geometry is constructed from which averaged integrated PDFs are generated (Section 4). These PDFs are independent of incoming directions and are used for all simulated events. An extended unbinned likelihood reconstruction is employed based on the direction and timing information of individual photons along with the per-module charge information, simulated via the Photon Propagation Code (Section 4) used by the IceCube Collaboration.
@@
We have derived the individual contributions from the major processes limiting the physics information, which subsequently impose limitations on the achievable resolutions. We have demonstrated (Fig. 9) that in-ice photon scattering and transverse momentum spread of shower particles (Section 3) are the dominant factors in limiting physics information. These contributions dominate over limitations arising from the directional resolution of PMTs in modules and module noise, both exhibiting minimal impacts on the reconstruction degradation. Due to negligible topological differences between CC, CC and NC events at these low energies, the contributions from resolution-limiting processes apply to all low-energy cascades observed in IceCube.
@@
The achievable reconstruction resolutions in ice are derived using integrated averaged PDFs and a toy simulation further indicates a 20% improvement in resolutions when angular and timing profiles of detected photons for every module are perfectly modeled instead of using an integrated distribution over all modules. These results are the first step towards estimating the potential improvement in the performance of reconstruction algorithms for GeV neutrino experiments, such as the IceCube Upgrade. Novel machine learning techniques offer improvements over traditional likelihood methods by better accommodating asymmetric detector geometries and thereby boosting reconstruction performance [Jessie_Micalf][GNN]. However, given the significant computational cost of training these algorithms, which increases with network complexity and benefits from larger training samples, the results from this study provide key insights into optimizing resource allocation to enhance low-energy reconstruction methods.
@@
Although the simulations and subsequent reconstruction in this paper have been performed for ice, the analysis is also applicable to water-based Cherenkov detectors like ORCA [KM3Net:2016zxf]. This is because seawater and ice are predominantly H20 leading to similar neutrino-nucleon cross-sections and subsequent particle propagation and overall shower profiles [ORCA_shower]. Therefore, it is expected that the limits on directional resolution by the shower spread will apply equally well in seawater. The proposed detector layout in ORCA has multi-PMT optical modules similar to mDOMs which are deployed on vertical strings. This detector architecture is expected to record photon distributions comparable to those in this study, though they will be shifted due to the different refractive index of water. Due to the lower scattering coefficient of water, photon scattering would have a smaller effect on limiting resolutions in water than in ice.